40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 848     Entry time: Mon Aug 18 17:37:14 2008
Author: rob 
Type: Update 
Category: Locking 
Subject: recovery progress 

I removed the beam block after the PSL periscope and opened the PSL shutter.

There was no MC Refl beam on the camera, so I decided to trust the PSL launch
and aligned the MC to the PSL beam. Here are the old and new values for
the MC angle biases:
 __Epics_Channel_Name______   __OLD_____    ___New___
 C1:SUS-MC1_PIT_COMM          4.490900        3.246900 
 C1:SUS-MC1_YAW_COMM          0.105500	      -0.912500
 C1:SUS-MC2_PIT_COMM          3.809700	      3.658600 
 C1:SUS-MC2_YAW_COMM          -1.837100	      -1.217100
 C1:SUS-MC3_PIT_COMM          -0.614200	      -0.812200
 C1:SUS-MC3_YAW_COMM          -3.696800	      -3.303800

After this, the beam looks a *little low* going into the Faraday Isolator.
Nonetheless, after turning on the IFO input steering PZTs, I was able to
quickly steer the PRM get a beam on the REFL camera and into the REFL OSA.
The PRM optical lever beam is also striking the quad.

I then used the ETMX optical lever as a reference for realigning. After
steering around the input PZTs and ITMX, I saw some flashes in Xarm trans, then got
it locked and ran the alignment script ~5 times. The arm power went
up to 0.9, so I tweaked the MC1 to put the MC refl beam back on MCWFS.
The XARM power then went up to .96. Good enough for now.

Then I started to try and re-align the YARM. Since the oplevs for both ITMY
and the BS are untrustworthy, I first tried to get the beam bouncing off ITMX
and the BS back into the AS OSA, to try and recover some BS alignment. This
didn't work, as the AS OSA may not be a good reference anyways. After
wandering around in the dark for a little while, I decided to try an automated
scan of the alignment space. I used the trianglewave script to scan
the angle biases of BS, ITMY, & ETMY, then looked at the trend of the transmitted
power to find the gps time when there were flashes. I then used
time_machine_conlog to restore the biases to that time. This was close
enough to easily recover the alignment. After several rounds of aligning &
centering oplevs, things look good.

Also locked a PRM. Will work on the DRM tomorrow.

I'm leaving the optics in their "aligned" states over night, so they can
start their "training."

Note: The MC is not staying locked. Needs investigation.

For tomorrow:

lock up the DRM
fix the mode cleaner
re-align mode cleaner to optimize beam through Faraday
re-align all optics again (will be much easier than today)
re-align beam onto all PDs after good alignment of suspended optics is established.
Attachment 1: flatlissa.png  19 kB  | Hide | Hide all
flatlissa.png
ELOG V3.1.3-